# Optimal Approximations for the Requirement Cut Problem on Sparse Graph Classes

### v2

@@ Abstract
We study the Requirement Cut problem, a generalization of numerous classical graph partitioning problems including Multicut, Multiway Cut, $k$-Cut, and Steiner Multicut among others. Given a graph with edge costs, terminal groups $S_{1},\ldots,S_{g}$ and integer requirements $r_{1},\ldots,r_{g}$; the goal is to compute a minimum-cost edge cut that separates each group $S_{i}$ into at least $r_{i}$ connected components. Despite many efforts, the best known approximation for Requirement Cut yields a double-logarithmic $\mathcal{O}(\log(g)\cdot\log(n))$ approximation ratio as it relies on embedding general graphs into trees and solving the tree instance.
In this paper, we explore two largely unstudied structural parameters in order to obtain single-logarithmic approximation ratios: (1) the number of minimal Steiner trees in the instance, which in particular is upper-bounded by the number of spanning trees of the graphs multiplied by $g$, and (2) the 0ptof series-parallel graphs. Specifically, we show that if the number of minimal Steiner trees is polynomial in $n$, then a simple LP-rounding algorithm yields an $\mathcal{O}(\log n)$-approximation, and if the graph is series-parallel with a constant depth then a refined analysis of a known probabilistic embedding yields a $\mathcal{O}(0pt\cdot\log g)$-approximation on series-parallel graphs of bounded depth. Both results extend the known class of graphs that have a single-logarithmic approximation ratio.
@@ 1 Introduction
Graph partitioning problems are a central topic in combinatorial optimization, with applications such as from network design [14], VLSI Layout [10], image segmentation [23, 4], parallel computing [16] and even bioinformatics [29]. In this paper, we study the Requirement Cut problem, a generalization of several well-known cut problems. Given an undirected graph $G=(V,E)$ with edge costs, a collection of $g$ terminal groups $\mathcal{S}=(S_{1},\dots,S_{g})\subseteq V$, and a requirement $r_{i}>1$ for each group, the goal is to compute a minimum-cost edge cut $C\subseteq E$ that separates each terminal group $S_{i}$ into at least $r_{i}$ components in $G^{\prime}=(V,E\setminus C).$
The Requirement Cut problem is a natural extension of multiple classical problems that have been studied for decades. For each bullet point, next to the name of the problem, we will reference a paper that appears to be the one that introduced the problem to the community. Inside each bullet points, we will reference papers and results that have the current best general and optionally other more specific parametrized approximation ratio or run-times:
@@
  • Multicut [27] : Given a set of terminal pairs $(s_{1},t_{1}),\dots,(s_{k},t_{k})$, the goal is to find a minimum-cost edge cut that separates each pair. This corresponds to the Requirement Cut problem when each terminal group consists of exactly two vertices and has a requirement of $r_{i}=2$. Garg et al. [21] gives the best known approximation for general graphs with a $\mathcal{O}(\log(k))$-approximation ratio via LP rounding (region growing). Friedrich et al. give a constant approximation for series-parallel graphs [19] and a more general [20] $\mathcal{O}(\log(tw))$ for graphs of treewidth $tw$.
@@
  • Multiway Cut [12]: Given a set of terminals $S=\{s_{1},\dots,s_{k}\}$, the objective is to find a minimum-cost edge cut that separates all terminals into distinct components. This is a special case of Requirement Cut where there is only one group and $r=k$. Chopra et al. [9] studied the linear program of this problem in depth before Calinescu et al. [11] improved the best approximation from $2-\frac{2}{k}$ to $1.5-\frac{1}{k}$. This problem also has seen a relatively new variant in the past few years called the Norm Multiway Cut[7, 6].
@@
  • Steiner-Multicut [31]: Given sets of terminals $\mathcal{S}=(S_{1},\dots,S_{k})$ each of size $(t_{1},\dots,t_{k})$, the goal is to find a minimum-cost edge cut that separates each set in at least 2 components. This corresponds to the Requirement Cut problem when each requirement is 2. This problem has particularly many applications has even been studied in a parametrized complexity setting [5]. Its best current approximation comes directly from the Requirement Cut and comes with a $\mathcal{O}(\log(k)\log(n))$ ratio. Moreover, it is a direct generalization of a less studied variant when $k=1$, the Steiner-Cut problem [28].
@@
  • Multi-Multiway Cut [2]: A generalization of Multiway Cut where multiple terminal groups must be separated into distinct connected components. This corresponds to the Requirement Cut problem when each terminal group has a requirement equal to its size. The best known approximation comes from the paper that introduced the problem [2] and is $\log(2k)$. Slightly later two more papers have studied the problem under different angles, [35] by adding label edges and [13] studying the problem on bounded branch-width graphs.
@@
  • k-Cut [22] : The problem of partitioning a graph into at least $k$ connected components using a minimum-cost edge cut. This can be viewed as a Requirement Cut instance where there is only $S_{1}=V$ as a terminal set and $r_{1}=k$. After the introduction of the problem that studied its parametrized complexity, Saran and Vazirani[34] found the current best approximation ratio $2-\frac{2}{k}$. More recently, Gupta et al. [24], gave tight bounds for the run-time of computing an optimal solution.
@@
  • Steiner k-Cut [8]: This is a generalization of the k-Cut problem, where $S_{1}\subseteq V,|S_{1}|\geq k$ and the objective is to find a minimum cost set of edges whose removal results in at least $r_{1}=k$ disconnected components, each containing a terminal. When $k=n$ this is exactly the k-Cut problem. The problem has been established and approximated best by Chekuri et al. [8] to a $2-\frac{2}{k}$ ratio. It is worth noting that their work has established the Linear Program formulation used to solve the Requirement Cut Problem introduced later.
Refer to caption[21][20][11][34][25][2][8][25](Section 3)(Section 5)
@@
These problems have been studied mostly through the lens of approximation algorithms. Some of them have received particular attention on specific classes of graphs with parameters such as treewidth, branchwidth or pathwidth.
The hardness of the Requirement Cut can be shown through any of these problems since they are all hard and are particular cases of it. A particular reduction that helps us with bounding the approximability on trees is a reduction from Set Cover taken from [32]. Consider a reduction from Set Cover to the Requirement Cut problem on trees. We construct a star graph where each leaf node corresponds to a set in the Set Cover instance, and the center node of the star connects to each of these leaves via an edge of cost $1$. For each element $j$ in the universe, we define a group consisting of the root and all leaf nodes whose corresponding sets contain element $j$. We assign a requirement of $2$ to each group. In this construction, selecting an edge in the tree (i.e., choosing a set in the Set Cover instance) ensures that the corresponding group is cut (i.e., its requirement is satisfied) if and only if that set covers one of the group’s elements. Hence, there is a one-to-one correspondence between feasible solutions of the Requirement Cut instance and valid set covers, and this correspondence preserves the total cost. Feige et al. [18] showed that Set Cover can not be approximated in polynomial time to a $\Omega(\log(n))$ ratio where $n$ is the size of the universe to cover. It follows that Requirement Cut on trees is hard to $\Omega(\log(g))$- approximate where $g$ is the number of groups.
Now that we have seen that the best approximation possible on trees for our problem is $\log(g)$, let us put the results of our paper in context. Note that, the reduction from Set Cover to Requirement Cut on trees works even when all the requirements are 2 (i.e. the Requirement Cut instance is a Steiner Multicut instance). In 1997, Klein, Plotkin, Rao and Tardos [31] presented an algorithm that yields $\mathcal{O}(\log^{3}(gt))$-approximation with $t=\max(t_{i})$ (the sizes of the sets) for the Steiner Multicut problem. This “triple log” approximation ratio remained the best for this problem until 2010. In 2010, the Requirement Cut problem was introduced by Nagarajan and Ravi [32] who proved a $\mathcal{O}(\log(n)\cdot\log(gR))$-approximation with $R=\max(r_{i})$. This approximation became, at the same time, the best approximation ratio for the Steiner Multicut problem. Their bound was later improved by the same set of authors along with Gupta [25] that got the approximation ratio down to $\mathcal{O}(\log(n)\cdot\log(g))$, effectively removing all dependency on the requirements. Both these approaches, that manage to reduce the power of “logs” in the approximation ratio from triple to double, involve very similar algorithmic steps:
@@
    Finding a good approximation algorithm for trees, $\log(gR)$ for [32] and $\log(g)$(tight) for [25].
@@
    Using Fakcharoenphol, Rao and Talwar (FRT) [17], they embed the graph into a tree and pay a $\log(n)$ factor for it.
@@
It is also worth noting that Gupta et al. [25] technically has a slightly better $\mathcal{O}(\log(k)\cdot\log(g))$-approximation with $k$ being the total number of terminals. They obtain this improvement through something that they call a slight generalization of FRT [17] in which they embed into a tree only the terminal vertices and hence only pay $\log(k)$ and not $\log(n)$. But this remains secondary in their paper since their main contribution was closing the gap between the $\log(gR)$ approximation on trees and the known lower bound of $\log(g)$ making it tight.
For the past few years, one goal of the community has been to beat those “double log” approximations with “single log” ones. Since we know that embedding into trees has a $\log(n)$ cost, (surprisingly) even for treewidth 2 graphs [26], and that a Requirement Cut problem instance on trees can not be approximated better than with a $\log(g)$ factor [18], we know that continuing this path of using tree embeddings will not result into a “single log” approximation. In this paper, we try something different over two different routes that we present in the next section.
@@ 1.1 Our Contributions
We address the following question: For which classes of graphs does the Requirement Cut problem admit optimal (single-log) approximation algorithms? To this end, we identify two structural parameters—(1) the number of spanning or minimal Steiner trees of the input graph and (2) the depth of series-parallel graphs — we show that both these relatively unstudied parameters offer better approximation for the Requirement Cut problem.
First, we show that for series-parallel graphs of constant 0pt, a refined analysis of Emek and Peleg’s [15] existing embeddings method yields an $\mathcal{O}(\log(g))$-approximation, extending the known $\mathcal{O}(log(g))$-approximations for trees to a broader class of graphs.
Second, we demonstrate that graphs with a polynomial number of spanning trees (or, more generally, number of minimal Steiner trees) admit an $\mathcal{O}(\log(n))$-approximation via a randomized LP-rounding scheme similar to the one given by Nagarajan and Ravi [32].
Our results broaden the landscape of instances that can be approximated with a “single logarithmic” factor and shed light on two unstudied parameters until now, $depth$ of series-parallel graphs and the order of magnitude of the number of spanning trees in the graph in the context of Requirement Cut.
@@
  • Instances with a Polynomial Number of Steiner Trees: A key parameter in our analysis of our Algorithm 1 of Section 4 depends on the instance. We will refer to it as $\sigma(G,\mathcal{S})$ for a given Requirement Cut instance. This function is equal to the number of distinct Minimal Steiner Trees (Steiner Trees for which all the leaves are terminal). This quantity is hard to compute [36] and so we will almost always upper-bound it by $\tau(G)\cdot g$ with $\tau(G)$ being the number of distinct spanning trees of the graph which is computable in polynomial time using Kirchhoff’s Matrix Tree Theorem initially stated in [30]. We will show that if $\tau(G)=\mathcal{O}(n^{c_{1}})$ then necessarily $\sigma(G,\mathcal{S})=\mathcal{O}(n^{c_{2}})$ when $c_{1}$ and $c_{2}$ are constants.
@@
  • Bounded-Depth Series-Parallel Graphs: Series-parallel graphs are well known for their recursive structure. They are often studied as a generalization of trees since they are, for example, representative of treewidth 2 graphs (trees have treewidth 1) and also are exactly $K_{4}$-minor-free graphs (trees are $K_{3}$-minor-free). In our paper, we introduce their depth as a key parameter. This parameter has never been studied to the best of our knowledge and we show a $\mathcal{O}(depth\cdot\log(g))$-approximation for the Requirement Cut problem on series-parallel graphs. Extending the class of graphs that have a $\mathcal{O}(\log(g))$-approximation to the series-parallel graph with any constant $depth$. This class of graph can additionally be seen as a generalization of melonic graphs which are used in other fields such as physics.
@@
In Section 3 we will give several classes of graphs and instances for which our results apply. Afterwards in Section 4 we give, prove and analyze our algorithm regarding graphs of polynomial number of Steiner trees. This comes with a proof about the Linear Program formulation of the Requirement Cut problem that is of its own interest (Lemma 4.1). Finally in Section 5, we refine Emek and Peleg’s [15] analysis to adapt it to the depth of series-parallel graphs. This approach provides an alternative to the known $\log(n)$ embeddings for series-parallel graphs.

### v3

@@ Abstract
We study the Requirement Cut problem, a generalization of numerous classical graph partitioning problems including Multicut, Multiway Cut, $k$-Cut, and Steiner Multicut among others. Given a graph with edge costs, terminal groups $S_{1},\ldots,S_{g}$ and integer requirements $r_{1},\ldots,r_{g}$; the goal is to compute a minimum-cost edge cut that separates each group $S_{i}$ into at least $r_{i}$ connected components. Despite many efforts, the best known approximation for Requirement Cut yields a double-logarithmic $\mathcal{O}(\log(g)\cdot\log(n))$ approximation ratio as it relies on embedding general graphs into trees and solving the tree instance.In this paper, we explore two largely unstudied structural parameters in order to obtain single-logarithmic approximation ratios: (1) the number of minimal Steiner trees in the instance, which in particular is upper-bounded by the number of spanning trees of the graphs multiplied by $g$, and (2) the 0ptof series-parallel graphs. Specifically, we show that if the number of minimal Steiner trees is polynomial in $n$, then a simple LP-rounding algorithm yields an $\mathcal{O}(\log n)$-approximation, and if the graph is series-parallel with a constant depth then a refined analysis of a known probabilistic embedding yields a $\mathcal{O}(0pt\cdot\log g)$-approximation on series-parallel graphs of bounded depth. Both results extend the known class of graphs that have a single-logarithmic approximation ratio.
@@ 1 Introduction
Graph partitioning problems are a central topic in combinatorial optimization, with applications such as from network design [donde_identification_2005], VLSI Layout [cong_multilevel_2003], image segmentation [grady_isoperimetric_2006, boykov_fast_2001], parallel computing [bader_graph_2013] and even bioinformatics [junker_analysis_2008]. In this paper, we study the Requirement Cut problem, a generalization of several well-known cut problems. Given an undirected graph $G=(V,E)$ with edge costs, a collection of $g$ terminal groups $\mathcal{S}=(S_{1},\dots,S_{g})\subseteq V$, and a requirement $r_{i}>1$ for each group, the goal is to compute a minimum-cost edge cut $C\subseteq E$ that separates each terminal group $S_{i}$ into at least $r_{i}$ components in $G^{\prime}=(V,E\setminus C).$The Requirement Cut problem is a natural extension of multiple classical problems that have been studied for decades. In the following, we present those problems, reference the paper that introduced it and summarize the current best general and parameterized approximation ratios:
@@
  • Multicut [hu_multi-commodity_1963] : Given a set of terminal pairs $(s_{1},t_{1}),\dots,(s_{k},t_{k})$, the goal is to find a minimum-cost edge cut that separates each pair. This corresponds to the Requirement Cut problem when each terminal group consists of exactly two vertices and has a requirement of $r_{i}=2$. Garg et al. [garg_approximate_1993] gives the best known approximation for general graphs with a $\mathcal{O}(\log(k))$-approximation ratio via LP rounding (region growing). Friedrich et al. give a constant approximation for series-parallel graphs [friedrich_primal-dual_2022] and a more general [friedrich_approximate_2023] $\mathcal{O}(\log(tw))$ for graphs of treewidth $tw$.
@@
  • Multiway Cut [dahlhaus_complexity_1994]: Given a set of terminals $S=\{s_{1},\dots,s_{k}\}$, the objective is to find a minimum-cost edge cut that separates all terminals into distinct components. This is a special case of Requirement Cut where there is only one group and $r=k$. Chopra et al. [chopra_multiway_1991] studied the linear program of this problem in depth and got a $2-\frac{2}{k}$ approximation ratio. After several improvements and new techniques used [calinescu_improved_2000, karger_rounding_2004, buchbinder_simplex_2018], this chain culmianted with a 1.2965 approximation by Sharma and Vondràk [sharma_multiway_2014] . This problem also has seen a relatively new variant in the past few years called the Norm Multiway Cut [chandrasekaran_l_p-norm_2021, carlson_approximation_2023] which is a variant where the goal is to minimize the norm of the edges in the boundaries of $k$-parts of the graph.
@@
  • Steiner-Multicut [klein_approximation_1997]: Given sets of terminals $\mathcal{S}=(S_{1},\dots,S_{k})$ each of size $(t_{1},\dots,t_{k})$, the goal is to find a minimum-cost edge cut that separates each set in at least 2 components. This corresponds to the Requirement Cut problem when each requirement is 2. This problem has particularly many applications has even been studied in a parametrized complexity setting [bringmann_parameterized_2016]. Its best current approximation comes directly from the Requirement Cut and comes with a $\mathcal{O}(\log(k)\log(n))$ ratio. Moreover, it is a direct generalization of a less studied variant when $k=1$, the Steiner-Cut problem [jue_near-linear_2019].
@@
  • Multi-Multiway Cut [avidor_multi-multiway_2007]: A generalization of Multiway Cut where multiple terminal groups must be separated into connected components. This corresponds to the Requirement Cut problem when each terminal group has a requirement equal to its size. The best known approximation comes from the paper that introduced the problem [avidor_multi-multiway_2007] and is $\log(2k)$. Slightly later two more papers have studied the problem under different angles, [shuguang_multi-multiway_2009] by adding label edges and [deng_multi-multiway_2013] studying the problem on bounded branch-width graphs.
@@
  • k-Cut [goldschmidt_polynomial_1994] : The problem of partitioning a graph into at least $k$ connected components using a minimum-cost edge cut. This can be viewed as a Requirement Cut instance where there is only one terminal set covering all the vertices and its requirement $r=k$ ($k$ parameter given by the problem). First, the 3-cut was introduced and solved in polynomial time by Hochbaum and Schmoys [hochbaum_ov_1985]. It was then extended to k-cut problem by Goldsmith and Hochbaum [goldschmidt_polynomial_1994] who gave a polynomial time algorithm for any fixed $k$. Later, by better understanding its parametrized complexity, Saran and Vazirani[saran_finding_1995] found the current best approximation ratio $2-\frac{2}{k}$. More recently, Gupta et al. [gupta_optimal_2021], gave tight bounds for the run-time of computing an optimal solution.
@@
  • Steiner k-Cut [chekuri_steiner_2006]: This is a generalization of the k-Cut problem, where $S_{1}\subseteq V,|S_{1}|\geq k$ and the objective is to find a minimum cost set of edges whose removal results in at least $r_{1}=k$ disconnected components, each containing a terminal. When $k=n$ this is exactly the k-Cut problem. The problem has been established and approximated best by Chekuri et al. [chekuri_steiner_2006] to a $2-\frac{2}{k}$ ratio. It is worth noting that their work has established the Linear Program formulation used to solve the Requirement Cut Problem introduced later.
Refer to caption[garg_approximate_1993][friedrich_approximate_2023][sharma_multiway_2014][saran_finding_1995][gupta_improved_2010][avidor_multi-multiway_2007][chekuri_steiner_2006][gupta_improved_2010](Section 3)(Section 5)
@@
These problems have been studied mostly through the lens of approximation algorithms. Some of them have received particular attention on specific classes of graphs with parameters such as treewidth, branchwidth or pathwidth.The hardness of the Requirement Cut can be shown through any of these problems since they are all hard and are particular cases of it. A particular reduction that helps us with bounding the approximability on trees is a reduction from Set Cover taken from [nagarajan_approximation_2010]. Consider a reduction from Set Cover to the Requirement Cut problem on trees. We construct a star graph where each leaf node corresponds to a set in the Set Cover instance, and the center node of the star connects to each of these leaves via an edge of cost $1$. For each element $j$ in the universe, we define a group consisting of the root and all leaf nodes whose corresponding sets contain element $j$. We assign a requirement of $2$ to each group.In this construction, selecting an edge in the tree (i.e., choosing a set in the Set Cover instance) ensures that the corresponding group is cut (i.e., its requirement is satisfied) if and only if that set covers one of the group’s elements. Hence, there is a one-to-one correspondence between feasible solutions of the Requirement Cut instance and valid set covers, which cost the same. Feige et al. [feige_threshold_1998] showed that Set Cover can not be approximated in polynomial time to a $\Omega(\log(n))$ ratio where $n$ is the size of the universe to cover. It follows that Requirement Cut on trees is hard to $\Omega(\log(g))$- approximate where $g$ is the number of groups.Now that we have seen that the best approximation possible on trees for our problem is $\log(g)$, let us put the results of our paper in context. Note that, the reduction from Set Cover to Requirement Cut on trees works even when all the requirements are 2 (i.e. the Requirement Cut instance is a Steiner Multicut instance). In 1997, Klein, Plotkin, Rao and Tardos [klein_approximation_1997] presented an algorithm that yields $\mathcal{O}(\log^{3}(gt))$-approximation with $t=\max(t_{i})$ (the sizes of the sets) for the Steiner Multicut problem. This “triple log” approximation ratio remained the best for this problem until 2010. In 2010, the Requirement Cut problem was introduced by Nagarajan and Ravi [nagarajan_approximation_2010] who proved a $\mathcal{O}(\log(n)\cdot\log(gR))$-approximation with $R=\max(r_{i})$. This approximation became, at the same time, the best approximation ratio for the Steiner Multicut problem. Their bound was later improved by the same set of authors along with Gupta [gupta_improved_2010] that got the approximation ratio down to $\mathcal{O}(\log(n)\cdot\log(g))$, effectively removing all dependency on the requirements. Both involve very similar algorithmic steps:
@@
    Finding a good approximation algorithm for trees, $\log(gR)$ for [nagarajan_approximation_2010] and $\log(g)$ for [gupta_improved_2010].
@@
    Using Fakcharoenphol, Rao and Talwar (FRT) [fakcharoenphol_tight_2004], they embed the graph into a tree and pay a $\log(n)$ factor for it.
@@
It is also worth noting that Gupta et al. [gupta_improved_2010] technically has a slightly better $\mathcal{O}(\log(k)\cdot\log(g))$-approximation with $k$ being the total number of terminals. They obtain this improvement through something that they call a slight generalization of FRT [fakcharoenphol_tight_2004] in which they embed into a tree only the terminal vertices and hence only pay $\log(k)$ and not $\log(n)$. But this remains secondary in their paper since their main contribution was closing the gap between the $\log(gR)$ approximation on trees and the known lower bound of $\log(g)$ making it tight.For the past few years, one goal of the community has been to beat those “double log” approximations with “single log” ones. Since we know that embedding into trees has a $\log(n)$ cost, (surprisingly) even for treewidth 2 graphs [gupta_cuts_2004], and that a Requirement Cut problem instance on trees can not be approximated better than with a $\log(g)$ factor [feige_threshold_1998], we know that continuing this path of using tree embeddings will not result into a “single log” approximation. In this paper, we try something different over two different routes that we present in the next section.
@@ 1.1 Our Contributions
We address the following question: For which classes of graphs does the Requirement Cut problem admit optimal (single-log) approximation algorithms? To this end, we identify two structural parameters—(1) the number of spanning or minimal Steiner trees of the input graph and (2) the depth of series-parallel graphs. We show that both these relatively unstudied parameters offer better approximation for the Requirement Cut problem.First, we demonstrate that graphs with a polynomial number of spanning trees (or, more generally, number of minimal Steiner trees) admit an $\mathcal{O}{\log(n)}$-approximation via a randomized LP-rounding scheme similar to the one given by Nagarajan and Ravi [nagarajan_approximation_2010].Second, we show that for series-parallel graphs of constant 0pt, a refined analysis of Emek and Peleg’s [emek_tight_2010] existing embeddings method yields an $\mathcal{O}{\log(g)}$-approximation, extending the known $\mathcal{O}{log(g)}$-approximations for trees to a broader class of graphs.The results broaden the landscape of instances that can be approximated with a “single logarithmic” factor and shed light on two unstudied parameters until now, depthdepth of series-parallel graphs and the order of magnitude of the number of spanning trees in a graph.
@@
  • Instances with a Polynomial Number of Steiner Trees: A key parameter in our analysis of Algorithm 1 of Section 4 depends on the instance. We refer to it as $\sigma(G,\mathcal{S})$ for a given Requirement Cut instance. This function is equal to the number of distinct Minimal Steiner Trees (Steiner Trees for which all the leaves are terminal). This quantity is hard to compute [valiant_complexity_1979] and so we will almost always upper-bound it by $\tau(G)\cdot g$ with $\tau(G)$ being the number of distinct spanning trees of the graph which is computable in polynomial time using Kirchhoff’s Matrix Tree Theorem initially stated in [kirchhoff_ueber_1847]. We show that if $\tau(G)=\mathcal{O}(n^{c_{1}})$ then necessarily $\sigma(G,\mathcal{S})=\mathcal{O}(n^{c_{2}})$ when $c_{1}$ and $c_{2}$ are constants.
@@
  • Bounded-Depth Series-Parallel Graphs: Series-parallel graphs are well known for their recursive structure. They are often studied as a generalization of trees since they are, for example, representative of treewidth 2 graphs (trees have treewidth 1) and also are exactly $K_{4}$-minor-free graphs (trees are $K_{3}$-minor-free). In our paper, we introduce their depth as a key parameter. This parameter has never been studied to the best of our knowledge and we show a $\mathcal{O}(depth\cdot\log(g))$-approximation for the Requirement Cut problem on series-parallel graphs. Extending the class of graphs that have a $\mathcal{O}(\log(g))$-approximation to the series-parallel graphs with any constant depthdepth. This class of graph can additionally be seen as a generalization of melonic graphs which are used in other fields such as physics.
@@
In Section 3 we will give several classes of graphs and instances for which our results apply. Afterwards in Section 4 we give, prove and analyze our algorithm regarding graphs of polynomial number of Steiner trees. This comes with a proof about the Linear Program formulation of the Requirement Cut problem that is of its own interest (Lemma 4.1). Finally in Section 5, we refine Emek and Peleg’s [emek_tight_2010] analysis to adapt it to the depth of series-parallel graphs. This approach provides an alternative to the known $\log(n)$ embeddings for series-parallel graphs. Due to space restrictions, some technical proofs are moved to the appendix; such statements are marked with a $\star$.